Solve the initial value probelm of y'-y=e^x, y(1)=0


1
Expert's answer
2021-06-28T03:36:55-0400

"\text{The integrating factor is } e^{-\int dx}=e^{-x}\\\text{Multiply the given differential equation by } e^{-x}
\\\implies e^{-x}y'-e^{-x}y=e^{-x}e^x\\\int d(e^{-x}y)=\int1dx\\e^{-x}y=x+c\\y=xe^x+ce^{x}\\\text{Recall that y(1)=0}\\\implies0=e^1 +ce^1\\\text{Therefore c = -1, hence the general solution is }y =xe^x-e^x"